OpenStudy (redman15):

The floor in a cube shaped room had a width of 2x + 5. If the area of the floor is 625 sq ft what is the value of "x"?

OpenStudy (sooobored):

cube - all sides are the same length

OpenStudy (royalranger):



OpenStudy (redman15):

Thx but it's multiple choicee A.310ft B.10ft C.15ft D.17.5ft

OpenStudy (troy121101):

then just plug each solution into the equation by substituting for x, and solve

OpenStudy (troy121101):

for example, instead of "2x - 5" you would write out, "2 * 10 - 5"

OpenStudy (troy121101):

or "2 * 310 - 5

OpenStudy (troy121101):

and then you should have your answer, if one of them equals 625ft^2, then that is the answer

OpenStudy (redman15):

So how do i kn wat my answer is if I just plugged in the answer choices

OpenStudy (troy121101):

sorry, I meant +5, not -5

OpenStudy (troy121101):

plug in your answer choices, and then solve the equation, if one of those solutions equals 625, then you have the answer to your question

OpenStudy (sooobored):

the area of a square is A=s^2 side length s = 2x+5 area is given as 625 ft^2 625=(2x+5)^2

OpenStudy (sooobored):

solve for x

OpenStudy (redman15):

Ok thx both of u I mean I thought y'all was gonna do it but I'll try

OpenStudy (troy121101):

I will help you out, but I wont give you the answer straight up